# Existence, uniqueness, and time-asymptotics of regular solutions in multidimensional thermoelasticity on domains with boundary

### v1

@@ Abstract
In the paper, we investigate the nonlinear thermoelasticity model in two- and three-dimensional convex and bounded domains. We propose new boundary conditions for the displacement. These conditions are not usual in thermoelasticity. Whereas, we posit the Neumann boundary condition for the temperature. We prove the existence of global, unique solutions for small initial data. The temperature positivity is also shown. Next, we investigate the long-time behavior of solutions. We show that the divergence-free part of the displacement oscillates. On the other hand, we prove that the potential part and the temperature are strongly coupled. The non-rotation part is strongly affected by heat propagation. It turns out that it tends to 0 as $x$ approaches infinity. Additionally, the temperature converges to a constant function.
@@ 1 Introduction
The constant $\nu\in\mathbb{R}$ is given. Constants $\lambda$ and $\mu$ are also fixed and are so-called the Lamé moduli (see [21, 28]). They satisfy
@@
where $\boldsymbol{\varepsilon}(\mathbf{u}):=\frac{1}{2}\left(\nabla\mathbf{u}+\nabla^{T}\mathbf{u}\right)$ is the linearized deformation tensor (see [30, 28]). In this form, we see a symmetry of the operator. However, for our purposes, the form in (1.2) is more suitable. We can treat the Lamé operator as a generalization of the Laplacian. It is more appropriate to consider the former than the latter from an engineering point of view.
We posit the Neumann boundary condition for the temperature. It is a natural setting in this context (see [11, 21]). It means that the heat cannot escape from the body. The standard boundary condition for the displacement is of the Dirichlet type (see [21, 32, 31, 20, 10, 11, 25, 26] and others). On the other hand, the Neumann boundary condition for $\mathbf{u}$ is also studied (see [21, 25, 29]) or even the Neumann condition for the stress tensor $\boldsymbol{\sigma}:=2\mu\boldsymbol{\varepsilon}(\mathbf{u})+\lambda\operatorname{div}\mathbf{u}I-\nu\theta I$. We and T. Cieślak considered the Dirichlet condition for $\mathbf{u}$ in one-dimensional domains in [4, 5]. We prove the existence and uniqueness of global strong solutions and investigate the time-asymptotic behavior. Our methods are hardly based on estimates, including the Fisher information. The application of them to the system with the Dirichlet condition for the displacement in higher dimensions presents certain technical challenges related to boundary integrals. Hence, first, we, in cooperation with T. Cieślak, M. Fuest, J. Lankiet, B. Muha, and S. Trifunović, studied the problem on the multidimensional torus in [6]. We omitted issues interlocked with the boundary conditions in this manner.
J. Březina and E. Feireisl study the compressible magnetohydrodynamics system in 3d in [8]. They propose a few physically grounded boundary conditions for the magnetic field vector. One of these is the conditions we have set for displacement, i.e.
@@
It is novel to posit (1.3) for the displacement in the context of thermoelasticity. Nevertheless, they fit well with our methods. They enable us to prove the same results as in [6]. Indeed, we are able to close calculations with a functional
@@
We refer the reader to [21, 11, 29] for a rigorous derivation of the model. Let us perceive that the system exhibits the conservation of energy. Indeed, we have the following proposition.
@@
In the second main result, we characterize the long-time behavior of solutions obtained in Theorem 1.4. The Helmholtz decomposition plays a crucial role here. It turns out that the potential part of the displacement behaves completely different than the divergence-free part as time approaches infinity. The latter demonstrates the independence of the former, as well as of the temperature. It will be shown that it satisfies the homogeneous wave equation and exhibits oscillatory behavior. We have a different issue with the curl-free part of the displacement. It is strongly coupled to the temperature. The heat propagation process causes the damping effect, which eventually leads to the vanishing of the rotation-free part of $\mathbf{u}$. On the other hand, the temperature tends to a constant function. An analogous asymptotic behavior we have in the case of (1.1) on the torus (see [6]). Let us also notice that the potential part of $\mathbf{u}$ and the temperature exhibit the same asymptotics as $u$ and $\theta$ in the 1d case (see [5]). Thus, we see that they behave as the second principle of thermodynamics predicts. We denote the rotation-free part of $\mathbf{u}$, $\mathbf{u}_{0}$ and $\mathbf{v}_{0}$ as $\boldsymbol{\chi}$, $\boldsymbol{\chi}_{0}$ and $\tilde{\boldsymbol{\chi}}$, respectively. Let us formulate our result regarding the long-time behavior of solutions of (1.1).
@@
Let us discuss the literature. As we mentioned above, problem (1.1) is being studied for the first time with the proposed boundary conditions for $\mathbf{u}$. However, the study of similar systems with different boundary conditions has a long history. First research was fairly devoted to studying linear thermoelasticity (see, for instance, [12, 23, 21] and references therein). One of the first approaches to the nonlinear models was paper [29]. The author proved the existence of local-in-time solutions for the one-dimensional nonlinear thermoelsticity. Moreover, the existence of global solutions has also been established for initial data near steady states. Further results concerning the local-in-time existence or the existence of global ones in a neighborhood of steady states with the nonnegative temperature were obtained in [25, 26, 27, 20].
@@
Until recently, there were no results concerning the global existence of solutions for nonlinear problems, even in one-dimensional cases. However, the authors got the existence and strong-weak uniqueness of global measure-valued solutions in [11]. We would like to emphasize that this result applies to higher dimensions. The global existence of unique solutions with a positive temperature in 1d was obtained in paper [4]. In article [6], thermoelasticity was studied on the multidimensional torus. The authors obtained the global unique solutions, as well as the positivity of temperature, for small initial data. Winkler approached the existence of weak solutions of the thermoviscoelasticity model in 1D in [31, 32]. The finite-time blow-up was demonstrated for a nonlinear problem in the 1d case in [13]. However, the authors studied the system with nonlinearity $(p(u_{x})u_{x})_{x}$ instead of the Laplacian (or the Lamé operator). The result was extended to a more general case (and a more physically appropriate one) in [19].
@@
Regarding the long-time behavior, we should mention works [29, 21, 10]. The asymptotics near steady states in a one-dimensional setting was studied there. The complete long-time behavior for the simplest nonlinear model for the one-dimensional setting was demonstrated in [5]. The full study of the time-asymptotics for higher dimensions is in [6]. However, the authors deal with the set without the boundary, i.e., the torus.
@@
At the end of the introduction, let us sketch the plan of the paper. We cite the general results which we shall use in the further part of the article in Section 2. In Subsection 2.1, we analyze basic facts about the entropy of (1.1). This result was already done in [11], but not with boundary conditions (1.3) for $\mathbf{u}$. Therefore, we decided to repeat the proof. Subsection 2.2 is devoted to the study of the Poisson equation with boundary conditions (1.3). It gives us the sequence of eigenfunctions of the Laplacian which satisfy (1.3). This, in turn, brings us the basis of $L^{2}(\Omega)$ and the basis of $\mathcal{H}$, which satisfies (1.3). The main burden of proof for the existence of solutions is found in Section 3. We prove a priori estimates there. The functional $\mathcal{F}$ is our main tool. We prove estimates that are independent of time for initial data such that $\mathcal{F}\leq D$. The constant $D$ occurs during the reasoning. To close the estimates, we require the $L^{\infty}$-estimate for the temperature. We prove it via the Moser iteration. Consequently, we obtain a bound for the temperature from below.
@@
Then, we proceed to the proof of the global existence for small data. We establish it in Section 4. First, in Subsection 4.1, we construct the sequence of approximate solutions. We employ the half-Galerkin method there. The same technique was applied in [11]. The earlier investigation of the eigenfunctions of the Laplacian with (1.3) enables the half-Galerkin construction. In Subsection 4.2, we prove the existence of global solutions of (1.1) for the initial values such that $\mathcal{F}\leq D$. The constant $D$ is taken from Section 3. The estimates in this section allow transition to the limit with the sequence of the approximate solutions. Then, it is shown that the limit functions are solutions. The final subsection, Subsection 4.3, is devoted to proving the uniqueness of solutions.
@@ 2 Preliminaries
This section is devoted to introducing general results that will be utilized later. We start by citing the lemma from [15]. It will be crucial in proving Lemma 3.1. It states that a certain integral, which will appear in the proof, is nonnegative.
@@
The next lemma presents the unusual inequality between the norms of the gradient of a function and its Laplacian. The proof can be found in [17].
@@
As in previous papers about thermoelasticity ([5, 6, 4]), the following inequality will also be important. It estimates the integral that will appear in the calculations. The proof can be found in [9].
@@
In the following theorem, the Helmholtz decomposition of a vector field is stated. It will be applied to decompose the displacement in Subsection 5.1. For the proof, see e.g. [16].
@@ 2.1 Entropy
for almost all $(t,x)\in(0,T)\times\Omega$. An analogous line of reasoning to that in [11] leads us to the following proposition.
@@ 2.2 Eigenfunctions of vector Laplacian
It is obvious that $B$ is bounded. Moreover, Theorem IV.4.8 in [7] and Remark 3.5 of Chapter I in [18] guarantee that $B$ is coercive on $\mathcal{H}$ with the norm $\|\cdot\|_{H^{1}(\Omega)}$. Hence, by the Lax-Milgram Theorem, we obtain that for any $\mathbf{f}\in L^{2}(\Omega)$ there exists a unique $\mathbf{w}\in\mathcal{H}$ such that
@@
Moreover, Theorem IV.9.6 of [7] gives us that $\mathbf{w}\in H^{2}(\Omega)$. Let us notice that there is a proof only for $d=3$, but it is analogous in the case when $d=2$. Therefore, we can state the following result.
@@
Now, we can apply the standard theory of symmetric and compact operators in Hilbert spaces (see [14, Theorem 7 of Appendix D]) to the operator solving problem (2.5). As an inference, we get the existence of an orthogonal basis (with $B$ defined in (2.4) as scalar product) $\{\boldsymbol{\varphi}_{k}\}$ of $\mathcal{H}$ consisting of eigenfunctions as in (2.3), which in addition is an orthonormal basis in $L^{2}(\Omega)$. Let us denote the eigenvalues of (2.3) by $\xi_{k}$. Because $\boldsymbol{\varphi}_{k}$ satisfy $\Delta\boldsymbol{\varphi}_{k}=\xi_{k}\boldsymbol{\varphi}_{k}$ in the interior of $\Omega$, due to the standard results for the regularity of solutions of elliptic equations (see [7, 14]), we get that $\boldsymbol{\varphi}_{k}\in C^{\infty}(\Omega)$. However, Theorem 2.6 says that $\boldsymbol{\varphi}_{k}\in H^{2}(\Omega)$ for all $k\in\mathbb{N}$. Albeit applying similar methods to those in [7], it can be shown that the eigenfunctions are arbitrarily well-regular as long as the boundary’s regularity allows it.
@@ 3 A priori estimates
It plays a crucial role in estimates. It consists of the Fisher information and higher-order derivatives of the displacement. Actually, this is a version of the functional that has been adopted to address our issue. Similar ones can be found in [4, 6, 5]. We start proving the inequality in the following lemma.
@@
Let us notice that the right-hand side of the inequality in the above theorem is finite because the bound from below of $\theta_{0}$ is positive. We will need a global estimate for the temperature in $L^{\infty}$ to complete our estimates. This will be inferred as a consequence of the following general lemma. We will prove it utilizing the Moser-Alikakos iteration (see [24, 1]). We already applied this method in the context of thermoelasticity in [5, 6]. However, it had been earlier used for 1d thermoelasticity in [13].
@@ 4 Existence and uniqueness of solutions
In this section, we prove the existence and uniqueness of solutions. First, we construct the sequence of approximate solutions. We utilize the half-Galerkin method there. Then, we apply the results from Section 3 to the sequence to obtain the estimates. This will enable us to reach the limit in the sequence. In this way, we will get the existence. Then, we will prove the uniqueness of solutions. A similar method to build the solution was applied in [6, 4, 11].
@@ 4.1 Approximate problem
We proceed in an analogous way as in [11, Proposition 1 and Lemma 4.2] to obtain the following result.
@@ 4.2 Existence of solutions
The standard theory of regularity for elliptic equations (see [14, 22, 2]) and the choice of an appropriate basis guarantee the sufficient regularity of $\mathbf{u}_{n}$ and $\theta_{n}$ to employ the results from Section 3. Inequalities (4.7) and (4.8) are immediate consequences of Theorem 3.6 and Theorem 3.7. Theorem 3.4, and (4.7) yield
@@
The right-hand side is bounded because of (4.2) and (4.7). Hence, we get (4.5). Moreover, from [7, Theorem III.4.3] we have
@@
Due to the Gelfand triple theorem (see, for instance, [7, 14]), we notice that
@@
Moreover, in virtue of the Aubin-Lions lemma (see [7]), we have

### v2

@@ Abstract
In the paper, we investigate the nonlinear thermoelasticity model in two- and three-dimensional convex and bounded domains. We propose new boundary conditions for the displacement. These conditions are not usual in thermoelasticity. Whereas, we posit the Neumann boundary condition for the temperature. We prove the existence of global, unique solutions for small initial data. The temperature positivity is also shown. Next, we investigate the long-time behavior of solutions. We show that the divergence-free part of the displacement oscillates. On the other hand, we prove that the potential part and the temperature are strongly coupled. The non-rotation part is heavily affected by heat propagation. It turns out that it tends to 0 as $x$ approaches infinity. Additionally, the temperature converges to a constant function.
@@ 1 Introduction
The constant $\nu\in\mathbb{R}$ is given. Constants $\lambda$ and $\mu$ are also fixed and are so-called the Lamé moduli (see Rackebook ; Slaughter ). They satisfy
@@
where $\boldsymbol{\varepsilon}(\mathbf{u}):=\frac{1}{2}\left(\nabla\mathbf{u}+\nabla^{T}\mathbf{u}\right)$ is the linearized deformation tensor (see Temamcont ; Slaughter ). In this form, we see a symmetry of the operator. However, for our purposes, the form in (1.2) is more suitable. We can treat the Lamé operator as a generalization of the Laplacian. It is more appropriate to consider the former than the latter from an engineering point of view.
We posit the Neumann boundary condition for the temperature. It is a natural setting in this context (see CMT ; Rackebook ). It means that the heat cannot escape from the body. The standard boundary condition for the displacement is of the Dirichlet type (see Rackebook ; Winkler1 ; Winkler2 ; Hrusa2 ; TC2 ; CMT ; Racke1 ; Racke2 and others). On the other hand, the Neumann boundary condition for $\mathbf{u}$ is also studied (see Rackebook ; Racke1 ; Slem ) or even the Neumann condition for the stress tensor $\boldsymbol{\sigma}:=2\mu\boldsymbol{\varepsilon}(\mathbf{u})+\lambda\operatorname{div}\mathbf{u}I-\nu\theta I$. We and T. Cieślak considered the Dirichlet condition for $\mathbf{u}$ in one-dimensional domains in BC ; BC2 . We prove the existence and uniqueness of global strong solutions and investigate the time-asymptotic behavior. Our methods are strongly based on estimates, including the Fisher information. The application of them to the system with the Dirichlet condition for the displacement in higher dimensions presents certain technical challenges related to boundary integrals. Hence, first, we, in cooperation with T. Cieślak, M. Fuest, J. Lankiet, B. Muha, and S. Trifunović, studied the problem on the multidimensional torus in Biest . We omitted issues interlocked with the boundary conditions in this manner.
J. Březina and E. Feireisl study the compressible magnetohydrodynamics system in 3d in fei . They propose a few physically grounded boundary conditions for the magnetic field vector. One of these is the conditions we have set for displacement, i.e.
@@
It is novel to posit (1.3) for the displacement in the context of thermoelasticity. Nevertheless, they fit well with our methods. They enable us to prove the same results as in Biest . Indeed, we are able to close calculations with a functional
@@
We refer the reader to Rackebook ; CMT ; Slem for a rigorous derivation of the model. Let us perceive that the system exhibits the conservation of energy. Indeed, we have the following proposition.
@@
In the second main result, we characterize the long-time behavior of solutions obtained in Theorem 1.4. The Helmholtz decomposition plays a crucial role here. It turns out that the potential part of the displacement behaves completely different than the divergence-free part as time approaches infinity. The latter demonstrates the independence of the former, as well as of the temperature. It will be shown that it satisfies the homogeneous wave equation and exhibits oscillatory behavior. We have a different issue with the curl-free part of the displacement. It is strongly coupled to the temperature. The heat propagation process causes the damping effect, which eventually leads to the vanishing of the rotation-free part of $\mathbf{u}$. On the other hand, the temperature tends to a constant function. An analogous asymptotic behavior we have in the case of (1.1) on the torus (see Biest ). Let us also notice that the potential part of $\mathbf{u}$ and the temperature exhibit the same asymptotics as $u$ and $\theta$ in the 1d case (see (BC2, )). Thus, we see that they behave as the second principle of thermodynamics predicts. We denote the rotation-free part of $\mathbf{u}$, $\mathbf{u}_{0}$ and $\mathbf{v}_{0}$ as $\boldsymbol{\chi}$, $\boldsymbol{\chi}_{0}$ and $\tilde{\boldsymbol{\chi}}$, respectively. Let us formulate our result regarding the long-time behavior of solutions of (1.1).
@@
Let us discuss the literature. As we mentioned above, problem (1.1) is being studied for the first time with the proposed boundary conditions for $\mathbf{u}$. However, the study of similar systems with different boundary conditions has a long history. First research was fairly devoted to studying linear thermoelasticity (see, for instance, lin1 ; lin2 ; Rackebook and references therein). One of the first approaches to the nonlinear models was paper Slem . The author proved the existence of local-in-time solutions for the one-dimensional nonlinear thermoelsticity. Moreover, the existence of global solutions has also been established for initial data near steady states. Further results concerning the local-in-time existence or the existence of global ones in a neighborhood of steady states with the nonnegative temperature were obtained in Racke1 ; Racke2 ; Racke3 ; Hrusa2 .
@@
Until recently, there were no results concerning the global existence of solutions for nonlinear problems, even in one-dimensional cases. However, the authors got the existence and strong-weak uniqueness of global measure-valued solutions in CMT . We would like to emphasize that this result applies to higher dimensions. The global existence of unique solutions with a positive temperature in 1d was obtained in paper BC . In article Biest , thermoelasticity was studied on the multidimensional torus. The authors obtained the global unique solutions, as well as the positivity of temperature, for small initial data. Winkler approached the existence of weak solutions of the thermoviscoelasticity model in 1D in Winkler2 ; Winkler1 . Moreover, thermoviscoelasticity in higher domains has been researched in Owczarek1 ; Owczarek . The finite-time blow-up was demonstrated for a nonlinear problem in the 1d case in DafHsi . However, the authors studied the system with nonlinearity $(p(u_{x})u_{x})_{x}$ instead of the Laplacian (or the Lamé operator). The result was extended to a more general case (and a more physically appropriate one) in Hrusa1 .
@@
Regarding the long-time behavior, we should mention works Slem ; Rackebook ; TC2 . The asymptotics near steady states in a one-dimensional setting was studied there. The complete long-time behavior for the simplest nonlinear model for the one-dimensional setting was demonstrated in BC2 . Moreover, the authors proved that the convergence to equilibrium is in exponential rate in thermorate . The full study of the time-asymptotics for higher dimensions is in Biest . However, the authors deal with the set without the boundary, i.e., the torus.
@@
At the end of the introduction, let us sketch the plan of the paper. We cite the general results which we shall use in the further part of the article in Section 2. In Subsection 2.1, we analyze basic facts about the entropy of (1.1). This result was already done in CMT , but not with boundary conditions (1.3) for $\mathbf{u}$. Therefore, we decided to repeat the proof. Subsection 2.2 is devoted to the study of the Poisson equation with boundary conditions (1.3). It gives us the sequence of eigenfunctions of the Laplacian which satisfy (1.3). This, in turn, brings us the basis of $L^{2}(\Omega)$ and the basis of $\mathcal{H}$, which satisfies (1.3). The main burden of proof for the existence of solutions is found in Section 3. We prove a priori estimates there. The functional $\mathcal{F}$ is our main tool. We prove estimates that are independent of time for initial data such that $\mathcal{F}\leq D$. The constant $D$ occurs during the reasoning. To close the estimates, we require the $L^{\infty}$-estimate for the temperature. We prove it via the Moser iteration. Consequently, we obtain a bound for the temperature from below.
@@
Then, we proceed to the proof of the global existence for small data. We establish it in Section 4. First, in Subsection 4.1, we construct the sequence of approximate solutions. We employ the half-Galerkin method there. The same technique was applied in CMT . The earlier investigation of the eigenfunctions of the Laplacian with (1.3) enables the half-Galerkin construction. In Subsection 4.2, we prove the existence of global solutions of (1.1) for the initial values such that $\mathcal{F}\leq D$. The constant $D$ is taken from Section 3. The estimates in this section allow transition to the limit with the sequence of the approximate solutions. Then, it is shown that the limit functions are solutions. The final subsection, Subsection 4.3, is devoted to proving the uniqueness of solutions.
@@ 2 Preliminaries
This section is devoted to introducing general results that will be utilized later. We start by citing the lemma from EvansEntropy . It will be crucial in proving Lemma 3.1. It states that a certain integral, which will appear in the proof, is nonnegative.
@@
The next lemma presents the unusual inequality between the norms of the gradient of a function and its Laplacian. The proof can be found in Fuestin .
@@
As in previous papers about thermoelasticity (BC2 ; Biest ; BC ), the following inequality will also be important. It estimates the integral that will appear in the calculations. The proof can be found in CFHS .
@@
In the following theorem, the Helmholtz decomposition of a vector field is stated. It will be applied to decompose the displacement in Subsection 5.1. For the proof, see e.g. FN .
@@ 2.1 Entropy
for almost all $(t,x)\in(0,T)\times\Omega$. An analogous line of reasoning to that in CMT leads us to the following proposition.
@@ 2.2 Eigenfunctions of vector Laplacian
It is obvious that $B$ is bounded. Moreover, Theorem IV.4.8 in Boyer and Remark 3.5 of Chapter I in G-R guarantee that $B$ is coercive on $\mathcal{H}$ with the norm $\|\cdot\|_{H^{1}(\Omega)}$. Hence, by the Lax-Milgram Theorem, we obtain that for any $\mathbf{f}\in L^{2}(\Omega)$ there exists a unique $\mathbf{w}\in\mathcal{H}$ such that
@@
Moreover, Theorem IV.9.6 of Boyer gives us that $\mathbf{w}\in H^{2}(\Omega)$. Let us notice that there is a proof only for $d=3$, but it is analogous in the case when $d=2$. Therefore, we can state the following result.
@@
Now, we can apply the standard theory of symmetric and compact operators in Hilbert spaces (see (Evans, , Theorem 7 of Appendix D)) to the operator solving problem (2.5). As an inference, we get the existence of an orthogonal basis (with $B$ defined in (2.4) as scalar product) $\{\boldsymbol{\varphi}_{k}\}$ of $\mathcal{H}$ consisting of eigenfunctions as in (2.3), which in addition is an orthonormal basis in $L^{2}(\Omega)$. Let us denote the eigenvalues of (2.3) by $\xi_{k}$. Because $\boldsymbol{\varphi}_{k}$ satisfy $\Delta\boldsymbol{\varphi}_{k}=\xi_{k}\boldsymbol{\varphi}_{k}$ in the interior of $\Omega$, due to the standard results for the regularity of solutions of elliptic equations (see Boyer ; Evans ), we get that $\boldsymbol{\varphi}_{k}\in C^{\infty}(\Omega)$. However, Theorem 2.6 says that $\boldsymbol{\varphi}_{k}\in H^{2}(\Omega)$ for all $k\in\mathbb{N}$. Albeit applying similar methods to those in Boyer , it can be shown that the eigenfunctions are arbitrarily well-regular as long as the boundary’s regularity allows it.
@@ 3 A priori estimates
It plays a crucial role in estimates. It consists of the Fisher information and higher-order derivatives of the displacement. Actually, this is a version of the functional that has been adopted to address our issue. Similar ones can be found in BC ; Biest ; BC2 . We start proving the inequality in the following lemma.
@@
Let us notice that the right-hand side of the inequality in the above theorem is finite because the bound from below of $\theta_{0}$ is positive. We will need a global estimate for the temperature in $L^{\infty}$ to complete our estimates. This will be inferred as a consequence of the following general lemma. We will prove it utilizing the Moser-Alikakos iteration (see Moser ; Ali ). We already applied this method in the context of thermoelasticity in BC2 ; Biest . However, it had been earlier used for 1d thermoelasticity in DafHsi .
@@ 4 Existence and uniqueness of solutions
In this section, we prove the existence and uniqueness of solutions. First, we construct the sequence of approximate solutions. We utilize the half-Galerkin method there. Then, we apply the results from Section 3 to the sequence to obtain the estimates. This will enable us to reach the limit in the sequence. In this way, we will get the existence. Then, we will prove the uniqueness of solutions. A similar method to build the solution was applied in Biest ; BC ; CMT .
@@ 4.1 Approximate problem
We proceed in an analogous way as in (CMT, , Proposition 1 and Lemma 4.2) to obtain the following result.
@@ 4.2 Existence of solutions
The standard theory of regularity for elliptic equations (see Evans ; Lady ; Amann ) and the choice of an appropriate basis guarantee the sufficient regularity of $\mathbf{u}_{n}$ and $\theta_{n}$ to employ the results from Section 3. Inequalities (4.7) and (4.8) are immediate consequences of Theorem 3.6 and Theorem 3.7. Theorem 3.4, and (4.7) yield
@@
The right-hand side is bounded because of (4.2) and (4.7). Hence, we get (4.5). Moreover, from (Boyer, , Theorem III.4.3) we have
@@
Due to the Gelfand triple theorem (see, for instance, Boyer ; Evans ), we notice that
@@
Moreover, in virtue of the Aubin-Lions lemma (see Boyer ), we have